OpenStudy (anonymous):

Identify the solution set of |x + 3| + 7 less than or equal to 10 and describe the graph in a complete sentence.

OpenStudy (jdoe0001):

\(\bf |x + 3| + 7 \le 10\implies |x + 3| \le 10-7\implies |x + 3| \le 3\\ \quad \\ |x + 3| \le 3\implies \begin{cases} +(x + 3) \le 3\\ \quad \\ \bf -(x + 3) \le 3 \end{cases}\)

OpenStudy (jdoe0001):

solve the inequalties, so you'd get 2 values for "x", 1 for each scenario as far as graphing them http://www.regentsprep.org/regents/math/algebra/ae85/grineqa.htm so you'd use a SOLID LINE, and you'd graph the x = SOME VALUE, line